OpenStudy (anonymous):

find the sum of the series. summation on n=1 going to infinity of (1/n(n+3))

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} 1/n(n+3)\]

OpenStudy (anonymous):

Break it up into partial fractions: \[\frac{1}{n(n+3)}=\frac{A}{n}+\frac{B}{n+3}\] Solve for A and B, then you have a series like this: \[\sum_{n=1}^\infty\left(\frac{A}{n}+\frac{B}{n+3}\right)\]